OpenStudy (zenmo):

Find an equation of the tangent line to the curve at the given point (Calculus Derivatives).

OpenStudy (zenmo):

Question #1

OpenStudy (freckles):

ok what did you find for y'?

OpenStudy (zenmo):

\[\lim_{h \rightarrow 0}\frac{ f(a+h)-f(a) }{ h }\] = \[\lim_{h \rightarrow 0}\frac{ [(3+h)^3-3(3+h)+1-(3^3-3(3)+1)] }{ h }\] . Putting more work on here second

OpenStudy (firekat97):

do you have to apply differentiation by first principles? :/

OpenStudy (zenmo):

\[(3+h)^3; (3+h)(3+h)=9+3h+3h+h^2=9+6h+h^2 ; (9+6h+h^2)(3+h) = 27+18h+3h^2+9h+6h^2+h^3\]

OpenStudy (anonymous):

Are you familiar with the power rule of differentiation? \[\frac{ d }{ dx }[x^n]=n*x^{n-1}\]

OpenStudy (freckles):

you might like pascal's triangle 4th row says 1 3 3 1 \[(3+h)^3=1(3)^3+3(3)^2h+3(3)h^2+h^3\]

OpenStudy (zenmo):

Our class isn't on there yet, this is the beginning of calculus for the end of Chapter 2

OpenStudy (freckles):

\[\lim_{h \rightarrow 0} \frac{3^3+3(3)^2h+3(3)h^2+h^3-3(3+h)+1-(3^3-3(3)+1)}{h}\]

OpenStudy (zenmo):

Sec. I might found the answer.

OpenStudy (zenmo):

I gotten 6x+1 as the answer.

OpenStudy (freckles):

ok let's go slow then you do see above the 3^3 's cancel right?

OpenStudy (freckles):

since 3^3-3^3=0 and also you should see the 3(3) also known as 9's cancel you know since -9+9=0

OpenStudy (zenmo):

This is how the book does it. Similar problem.

OpenStudy (freckles):

you should also see the 1's cancel since 1-1=0

OpenStudy (freckles):

\[\lim_{h \rightarrow 0} \frac{3^3+3(3)^2h+3(3)h^2+h^3-3(3+h)+1-(3^3-3(3)+1)}{h} \\ \lim_{h \rightarrow 0} \frac{3(3)^2h+3(3)h^2+h^3-3h}{h}\]

OpenStudy (freckles):

now divide top and bottom by h

OpenStudy (zenmo):

\[\lim_{h \rightarrow 0}(h^2+9h+27)\] Did u get that?

OpenStudy (freckles):

should wind up with \[\lim_{h \rightarrow 0} (3(3)^2+3(3)h+h^2-3)\]

OpenStudy (freckles):

something happened to your -3 part above

OpenStudy (freckles):

should be \[\lim_{h \rightarrow 0} (h^2+9h+27-3)\]

OpenStudy (freckles):

so your slope on your homework is off which will also make your y-intercept you found off too

OpenStudy (freckles):

please let me know if I need to redo the algebra for you if you don't understand

OpenStudy (freckles):

I guess you can call it algebra/arithmetic whatever that stuff is

OpenStudy (zenmo):

\[\lim_{h \rightarrow 0}(h^2+9h-24)\] that correct? I'm looking at calculations at the moment.

OpenStudy (freckles):

well 27-3 is 24 not -24 :p

OpenStudy (zenmo):

By using the point (3,19) and using the point-slope form for slope M=24. I get \[y=24x+19\]

OpenStudy (freckles):

y-19=f'(3)(x-3) y-19=24(x-3) y-19=24x-72 hmmm...-72+19...

OpenStudy (zenmo):

\[y=24x+53\]

OpenStudy (freckles):

I hate to be that guy but -72+19 is -53

OpenStudy (zenmo):

Yea so the answer is \[y=24x-53\] ?

OpenStudy (freckles):

yep

OpenStudy (freckles):

do you understand how we got the slope to be 24?

OpenStudy (zenmo):

Yea, turns out I was messing up on simple arithmetic/algebra breaks for: adding/subtracting/factoring etc.

OpenStudy (zenmo):

Yea I understood the core concepts, just messed up with the things from above.

OpenStudy (zenmo):

Time for a study break haha

OpenStudy (zenmo):

Thanks a lot for sticking around and helping. :)

OpenStudy (freckles):

we all mess on arithmetic just go pet a cat and then come back you will be already (that is my only study advice; cats for the win!)